# Coherent manipulation of interacting electron qubits on solid neon

### v4

@@
(August 24, 2025)
Solid neon has recently emerged as a pristine material host for electron qubits. Single electron-on-solid-neon (eNe) charge qubits have shown extraordinarily long coherence times and high operation fidelities. Realizing two-qubit gates in this platform is the next major step for practical quantum information processing. In this work, we demonstrate frequency- and time-domain coherent manipulation of multiple eNe charge qubits that are coupled by charge-charge interactions. Cross-resonance and bSWAP two-qubit gates are implemented, laying the foundation for universal quantum computing. Inter-qubit coupling strength exceeding 60 MHz is observed, promising fast and high-fidelity two-qubit gates. These results highlight the potential to develop the eNe qubit platform into a compelling quantum computing architecture.
Electrons on solid neon (eNe) is an emerging solid-state qubit platform [1, 2, 3, 4]. Owing to the ultraclean environment provided by the noble-element neon, qubits based on the motional (charge) states of eNe have shown exceedingly long coherence [2]. The inherent electrical addressability of charge states allows for straightforward control and readout by microwave photons in a circuit quantum electrodynamics architecture (cQED) [5, 6], significantly simplifying the circuit design. The small footprint of eNe charge qubits offers another unique advantage for high-volume integration within quantum circuits. To further advance the eNe platform towards a useful quantum information architecture, it is essential to construct systems with multiple interacting qubits and achieve coherent manipulation of all these qubits [7, 4]. In our previous work, we observed two eNe charge qubits trapped around the same designed trapping area and coupled through microwave photons in a common superconducting resonator [2]. Later theoretical and experimental studies show that an imperfect neon surface may play an important role in co-trapping multiple electrons [8, 9, 10]. However, all our observed through-resonator inter-qubit couplings so far are not strong enough for coherent real-time two-qubit gate operations.
@@
In this work, we achieve the first coherent manipulation of multiple interacting electron qubits on solid neon. The rough neon surface provides a natural mechanism to pack several electrons at close distances with strong charge-charge interactions surpassing the qubits’ decoherence rates [11, 4]. In a two-qubit system, spectroscopic measurements show an inter-qubit coupling strength of $J/2\pi=$ 3.35 MHz. By driving one of the two qubits, which is strongly coupled with the superconducting resonator, we are able to demonstrate the cross-resonance and bSWAP types of all-microwave two-qubit gates [12]. The experimental results match well the theoretical simulations based on measured parameters. Furthermore, we observe even stronger coupling in a three-qubit system, with a maximum inter-qubit coupling strength of $J/2\pi=62.5$ MHz. These results lay the foundation for scalable quantum information processing with this emerging platform. The local charge-charge coupling between qubits on solid neon may also provide a unique testbed for quantum simulation of correlated many-electron systems in an ultrapure near-vacuum environment [13, 14].
Multi-qubit systems
Fig. 1. : Schematic of multi-qubit coupling on neon-covered superconducting resonator. (A) Two closely arranged qubits trapped at the open end of a splitting co-planar waveguide superconducting resonator (pink), surrounded by electrical gates (yellow). The blue and gray structures are the intrinsic silicon substrate and ground, respectively. In the experiment, varying gate voltage is symmetrically applied on the left and right trap guards ($\Delta V_{\text{tg}}$) or on the resonator pins ($\Delta V_{\text{res}}$) to tune the qubits’ transition frequency. (B) Schematic of the interactions between a two-qubit system and the superconducting resonator, where a bright qubit $\mathrm{Q_{b}}$ and a dark qubit $\mathrm{Q_{d}}$ are coupled with strength $J$ and $\mathrm{Q_{b}}$ is coupled to the resonator with strength $g_{\mathrm{b}}$. Due to the orthogonal dipole alignment to the field, $\mathrm{Q_{d}}$ is invisible to the resonator. (C) Energy level schematic of the two-qubit system showing possible transitions.
Multiple electron qubits can be trapped in the same trapping area on our device, as shown in Fig. Coherent manipulation of interacting electron qubits on solid neonA. They can be driven and read out through microwave photons in the resonator. The electron-photon coupling strength is determined by the magnitude and alignment between the electric dipole moments of the qubits and the local electric field of the microwave photons. For two closely arranged qubits, $\mathrm{Q}_{i}$ and $\mathrm{Q}_{j}$, they can couple through not only microwave photons but also direct charge-charge interactions. The system Hamiltonian can be written as
$\hat{H}_{\mathrm{sys}}=\hat{H}_{\mathrm{r}}+\hat{H}_{\mathrm{q}}+\hat{H}_{\mathrm{r\text{-}q}}+\hat{H}_{\mathrm{q\text{-}q}},$ (1)

### v5

@@
Electrons trapped on solid neon surfaces serve as low-noise charge qubits with long coherence times and high operational fidelities. Such charge qubits offer full electrical control and compact device footprints, convenient for scaling up with quantum circuits. Realizing two-qubit gates on this platform is a critical step towards practical quantum information processing. In this work, we report the first experimental demonstration of coherent manipulation of multiple interacting electron-on-solid-neon (eNe) charge qubits. By exploiting the electrons naturally confined in close proximity by the surface structures of solid neon, we have achieved a direct qubit-qubit coupling strength of up to 62.5 MHz, as well as implemented cross-resonance (CR) and bSWAP two-qubit gates using global microwave drives. The natural electron confinement by solid neon mitigates the high-density-wiring challenge, simplifies the multi-qubit control, and establishes a unique path to scale up the eNe qubit platform.
Scalable quantum processors require the constituent qubits to have long coherence, fast and accurate gates, and high-connectivity integration. To that extent, electron-based qubits have unique advantages for large-volume solid-state system scaling, thanks to electrons’ small footprints. For example, spin qubits based on electrons in gate-defined quantum dots or implanted dopants exhibit very long coherence, and form multi-qubit arrays [burkard2023semiconductor, Petta2005coherent, borsoi2024shared, vandersypen2017interfacing, zwerver2022qubits, hu2025single]. However, this scaling requires complex gate stacks for electron confinement and spin detection at the individual qubit scale. Meanwhile, the necessity for global or local magnetic field gradients to address electron spin adds another layer of hardware complexity for integration. Those overheads impede the volume growth of electron-based qubit arrays [vandersypen2017interfacing, borsoi2024shared].
Electron-on-solid-neon (eNe) offers an emerging solid-state qubit platform that combines excellent coherence and minimal device overhead [zhou2022single, zhou2024electron, guo2024quantum, jennings2024quantum]. In this system, single electrons sit in a near-vacuum environment a few nanometers above the neon surface, where the noble-gas substrate provides superior noise isolation [li2025noise]. Qubits based on the charge (motional) states of eNe have shown exceedingly long coherence [zhou2024electron]. The inherent electrical addressability of charge states allows for straightforward and high-fidelity single-qubit gates with microwave photons in a circuit quantum electrodynamics (cQED) architecture [schuster2010proposal, blais2021circuit], significantly reducing the hardware complexity compared to electron spin qubits.
To further advance the eNe platform towards a functional quantum information architecture, it is essential to construct systems with multiple interacting qubits and realize their coherent manipulation [platzman1999quantum, jennings2024quantum]. Previously, we observed two eNe charge qubits co-trapped on the same device and simultaneously coupled to a common superconducting resonator [zhou2024electron]. However, the through-resonator inter-qubit coupling was not strong enough for coherent real-time two-qubit gate operations. Nonetheless, recent theoretical and experimental studies suggest that solid-neon surface morphology offers a natural mechanism to trap multiple electrons at close distances [kanai2024single, li2025noise, zheng2025surface]. It shows the promise of creating strong interactions between adjacent qubits maintained solely by the neon surface. This can greatly simplify the originally complex electrodes in typical electron-based qubit systems [vandersypen2017interfacing] and reduce gate-introduced electrical noise.
Fig. 1. : Schematic of multi-qubit coupling on solid neon. a, Solid neon creates multi-qubit host spots and supports inter-qubit coupling at a short distance. b, Schematic of the interactions between a two-qubit system on solid neon and a superconducting resonator, where a bright qubit $\mathrm{Q_{b}}$ and a dark qubit $\mathrm{Q_{d}}$ are coupled with strength $J$, and $\mathrm{Q_{b}}$ is coupled to the resonator with strength $g_{\mathrm{b}}$. Due to the orthogonal dipole alignment to the field, as marked by the black arrow, $\mathrm{Q_{d}}$ is invisible to the resonator. c, Energy-level schematic of the two-qubit system in b, illustrating the possible transitions. Both the cross-resonance (CR) and the bSWAP type two-qubit operations are shown, where the $\mathrm{Q_{d}}$ are driven by microwave gates mediated through the $\mathrm{Q_{b}}$.
@@
In this work, we explore this novel potential offered by solid neon and achieve the first coherent manipulation of multiple interacting electron qubits on this platform. We utilize the neon surface structure as a natural mechanism to pack several electrons at close distances to establish strong charge-charge interactions [beysengulov2024coulomb, jennings2024quantum]. Further, we prove that the natural confinement of electrons by solid neon offers a unique engineering resource to qubit-resonator and inter-qubit coupling, enabling multi-qubit operation with simplified wiring. In a two-qubit system, spectroscopic measurements show an inter-qubit coupling strength of $J/2\pi$ = 3.35 MHz, surpassing the qubits’ decoherence rates. By driving one of the two qubits, which is the only one strongly coupled with the superconducting resonator, we demonstrate the cross-resonance and bSWAP types of all-microwave two-qubit operations [krantz2019quantum], with just a single direct-current (DC) qubit bias. The experimental results match well with the theoretical simulations based on the measured parameters. Furthermore, we observe even stronger coupling in a three-qubit system, with a maximum qubit interaction strength of $J/2\pi$ = 62.5 MHz. Together with previously achieved long coherence and noise isolation, this result on coherent multi-qubit manipulation completes the eNe platform with crucial components to overcome key scaling challenges faced by electron-based qubits.
Multi-qubits on solid neon
The solid neon film can support the trapping of multiple electrons within a compact area without introducing complex electrical gate stacks at the individual qubit scale. Experimental and theoretical investigations show that surface structures at the 10-nm scale facilitate electron trapping on solid neon [zheng2025surface, kanai2024single]. Those surface structures can originate from the transfer of the underneath substrate roughness or the solidification process during the neon layer growth [leiderer2025surface]. As illustrated in Fig. Coherent manipulation of interacting electron qubits on solid neona, the solid neon film naturally supports closely spaced electron traps and enables local qubit coupling through strong charge-charge interactions.
The trapping of electrons on solid neon films also creates the possibilities for interaction engineering with simplified circuit wiring. In the cQED architecture, electrons can be driven and read out through microwave photons in the resonator. The electron-photon coupling strength is determined by the magnitude of the qubit’s dipole moment and its alignment with the local microwave (MW) electric field. The neon surface morphology can lead to distinct electron-photon couplings [koolstra2019coupling, schuster2010proposal]. As depicted in Fig. Coherent manipulation of interacting electron qubits on solid neonb, inside a closely spaced two-qubit coupled system, one of the qubits with a dipole orthogonal to the MW field of a superconducting resonator exhibits a near-zero qubit-resonator coupling strength and negligible resonator dispersive shift. We refer to this qubit as the “dark” qubit $\mathrm{Q_{d}}$. The other qubit, which has a dipole significantly aligned to the MW field, is coupled to the resonator with a strength of $g_{\mathrm{b}}$. This qubit is referred to as the “bright” qubit $\mathrm{Q_{b}}$ and its quantum states can be read out dispersively. The charge-charge interaction between the two qubits results in the inter-qubit coupling strength $J$.
Fig. 2. : Spectroscopic characterization of coherent two-qubit coupling on solid neon. a, Continuous-wave (CW) two-tone qubit spectroscopy with avoided crossings on the resonator’s phase response. The dashed lines indicate the bare transition frequencies of the dark qubit $\mathrm{Q_{d}}$ and the bright qubit $\mathrm{Q_{b}}$. b, The line cut at the avoided crossing shows a splitting corresponding to a charge-charge interaction strength of $J/2\pi=3.35$ MHz, which exceeds the hybrid linewidth $(\gamma_{\textrm{b}}+\gamma_{\textrm{d}})/4\pi=0.56$ MHz. c, Qubits spectroscopy measured with high-power pulsed readout, in which $\mathrm{Q_{b}}$ served as the mediator between $\mathrm{Q_{d}}$ and the resonator. The inset shows the pulse sequence: a long square pulse (10 µs) with variable frequency $f_{\mathrm{dr}}$ and power $P_{\mathrm{dr}}$ drive the system to a specific state, followed by a short square probe pulse (0.7 µs) at the resonator frequency $f_{\mathrm{r}}$ with approximately 120-120 dBm power reaching the resonator. The readout pulse induces the frequency crossing of the two qubits and the swap of their populations, revealing the state information of the dark qubit $\mathrm{Q_{d}}$. Drive pulse power $P_{\text{dr}}$ reaching the resonator’s input coupler increases from 73-73 dBm to 63-63 dBm, activating transitions corresponding to the CR ($|0_{\mathrm{b}}0_{\mathrm{d}}\rangle\rightarrow|0_{\mathrm{b}}1_{\mathrm{d}}\rangle$) and bSWAP ($|0_{\mathrm{b}}0_{\mathrm{d}}\rangle\rightarrow|1_{\mathrm{b}}1_{\mathrm{d}}\rangle$) two-qubit operations. The measurements were taken when the system was biased at $\Delta V_{\rm tg}=\,$0 V. The phase curves are off-set for visualization.
The structurally induced dipole orientations distribution enables coherent multi-qubit operations to be driven with only global control, reducing the complexity of the local gate stacks. The two-qubit system described above supports all-microwave-driven two-qubit operations, including cross-resonance (CR) and bSWAP gates, as presented in Fig. Coherent manipulation of interacting electron qubits on solid neonc. All-microwave gates are widely used for fixed-frequency qubits thanks to their easy implementation, which also preserves qubit coherence by avoiding electrostatic and flux bias [krantz2019quantum]. For the CR gate, $\mathrm{Q_{d}}$ is excited by driving $\mathrm{Q_{b}}$ at the frequency of $\mathrm{Q_{d}}$, generating the transition $|0_{\mathrm{b}}0_{\mathrm{d}}\rangle\rightarrow|0_{\mathrm{b}}1_{\mathrm{d}}\rangle$ (refs. [rigetti2010fully, magesan2020effective, chow2011simple]). For the bSWAP gate, $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ are driven monochromatically through a two-photon process near the frequency middle point between the two, generating the transition $|0_{\mathrm{b}}0_{\mathrm{d}}\rangle\rightarrow|1_{\mathrm{b}}1_{\mathrm{d}}\rangle$ (refs. [poletto2012entanglement, roth2017analysis]).
Two-qubit spectroscopy
We have experimentally achieved strong charge-charge coupling between eNe qubits, assisted by the neon surface, as proposed. See Methods for the details of the superconducting microwave resonator, neon growth and electrons deposition. A two-qubit system was probed by the continuous wave (CW) two-tone qubit spectroscopy while tuning one electrode gate ($\Delta V_{\rm tg}$, see Methods). Two clear avoided crossings which indicate the qubit-qubit strong coupling appear in the spectrum in Fig. Coherent manipulation of interacting electron qubits on solid neona. One of the qubits is visible in the CW spectroscopy and referred to as the “bright” qubit $\mathrm{Q_{b}}$, while the other non-detectable one is denoted as the “dark” qubit $\mathrm{Q_{d}}$. The magnitude of the splitting represents an inter-qubit coupling strength of $J/2\pi$ = 3.35 MHz, exceeding the linewidth of the hybrid states $(\gamma_{\textrm{b}}+\gamma_{\textrm{d}})/4\pi$ = 0.56 MHz, as shown in Fig. Coherent manipulation of interacting electron qubits on solid neonb.
Individually, $\mathrm{Q_{b}}$ has a hyperbolic frequency dependence on $\Delta V_{\rm tg}$, with a charge sweet spot (SS) near 5.71 GHz, which is about 40 MHz above the resonator frequency. We estimate the coupling strength between $\mathrm{Q_{b}}$ and resonator to be $g_{\mathrm{b}}/2\pi$ = 3.76 MHz, based on the dispersive pushing at the qubit’s SS (Extended Data Fig. 1).
Fig. 3. : Time-domain all-microwave two-qubit operations on solid neon. a and c, Drive pulse length-frequency (a) and drive pulse amplitude-frequency (c) Rabi measurements, following the pulse sequences depicted in the inset. In (a), the drive pulse power is fixed at 70-70 dBm, with varying length from 0 to 2.0 µs. In (c), the drive pulse length is fixed at 0.8 µs, with output amplitude varying from 0.0 V to 0.25 V, corresponding to zero power to 57-57 dBm reaching the resonator input coupler. The parameters of the readout pulses are the same as described in the Two-qubit Spectroscopy section. The system is biased with a single DC gate at $\Delta V_{\mathrm{tg}}=0$. Measurement results show the two oscillations corresponding to cross-resonance (CR, purple arrows) and bSWAP (magenta arrows) two-qubit gates. b and d, Numerically simulated excited state population of $\mathrm{Q_{b}}$ after the readout pulses, corresponding to the Rabi measurements in (b) and (d). The CR and bSWAP oscillation features are reproduced with experimentally measured parameters from the spectroscopy and decoherence characterization, see Methods for details.
In contrast, the direct interaction between $\mathrm{Q_{d}}$ and the resonator is not detectable. Meanwhile, the secondary frequency shift of the resonator caused by the dispersive push of $\mathrm{Q_{d}}$ on $\mathrm{Q_{b}}$ is also small. Together, these lead to a negligible $\mathrm{Q_{d}}$-state-dependent resonator response and the absence of visible $\mathrm{Q_{d}}$ spectrum in Fig. Coherent manipulation of interacting electron qubits on solid neona. In addition, the spectroscopic symmetry of the two avoided crossings in against $\Delta V_{\text{tg}}$ indicates that the frequency of the dark qubit $\mathrm{Q_{d}}$ is much less sensitive to $\Delta V_{\rm tg}$ compared to that of $\mathrm{Q_{b}}$, remaining nearly constant at 5.726 GHz over the voltage range scanned in Fig. Coherent manipulation of interacting electron qubits on solid neona. These observations reveal that the $\mathrm{Q_{d}}$’s dipole is nearly orthogonal to both the MW and DC gate fields, mainly defined by the neon-induced structural confinement.
We attribute the observed two-qubit coupling to charge-charge interactions at short distances. Due to the distinct electron-photon coupling strengths of the two qubits individually with the resonator, we can rule out the case of virtual microwave photon exchange between the two qubits through the resonator [majer2007coupling].
Although we cannot readout the dark qubit through the resonator directly, bringing $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ on resonance allows their populations to swap, which reveals the state of $\mathrm{Q_{d}}$ prior to that. This resonance condition is achieved via the ac-Stark shift induced by the readout pulse. Once resonant, the states exchange, and a subsequent measurement of $\mathrm{Q_{b}}$ reveals the information of the original $\mathrm{Q_{d}}$ state. A weak pulse provides an insufficient shift, preventing the swap and leaving $\mathrm{Q_{d}}$ unmeasured, as in the CW two-qubit spectroscopy of Fig. Coherent manipulation of interacting electron qubits on solid neona. We simulated the pulsed readout process with varied probe power, as detailed in Methods, to illustrate the state swap dynamics during the 0.7 µs readout pulse used in our experiments (Extended Data Fig. 4).
With this method, we perform pulsed qubit spectroscopy when biased at $\mathrm{Q_{b}}$’s SS, as shown in Fig. Coherent manipulation of interacting electron qubits on solid neonc. At low drive power of 73-73 dBm reaching the resonator, we observe two transitions at 5.711 GHz and 5.726 GHz, corresponding to the frequencies of $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$. The appearance of $|0_{\mathrm{b}}0_{\mathrm{d}}\rangle\rightarrow|0_{\mathrm{b}}1_{\mathrm{d}}\rangle$ transition indicates that the long drive pulse pumps $\mathrm{Q_{d}}$ to a certain population, corresponding to the CR transition of a two-qubit coupled system [chow2011simple, rigetti2010fully], as depicted in Fig. Coherent manipulation of interacting electron qubits on solid neonc. With increasing drive power, an additional transition peak ($|0_{\mathrm{b}}0_{\mathrm{d}}\rangle\rightarrow|1_{\mathrm{b}}1_{\mathrm{d}}\rangle$) emerges near the midpoint between the $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ frequencies, located at 5.718 GHz. This corresponds to the bSWAP type of two-qubit operation in Fig. Coherent manipulation of interacting electron qubits on solid neonc, involving two-photon excitation [poletto2012entanglement, roth2017analysis]. The fact that $2\omega_{|1_{\mathrm{b}}1_{\mathrm{d}}\rangle}\simeq\omega_{|0_{\mathrm{b}}1_{\mathrm{d}}\rangle}+\omega_{|1_{\mathrm{b}}0_{\mathrm{d}}\rangle}$ suggests the effective longitudinal (ZZ) interaction in the system is negligible [beysengulov2024coulomb, fors2024comprehensive]. In summary, through the frequency-domain characterization, we have observed the coherent coupling between two qubits on solid neon, which can support both the CR and the bSWAP types of two-qubit operation.
Two-qubit gate operations
Fig. 4. : Spectroscopic characterization of a three-qubit coupled system on solid neon. a, Two-tone qubit spectroscopy reveals a three-qubit coupled system, with the inter-qubit coupling strengths $J_{\mathrm{12}}$ and $J_{\mathrm{23}}$ shown in the inset. Via $\mathrm{Q}_{1}$, the system is probed at the resonator frequency while being driven with a second tone at varied frequencies. The dashed line marks the avoided crossing as in (b). b, The line cut shows the large charge-charge interaction strength between $\mathrm{Q}_{1}$ and $\mathrm{Q}_{2}$ of $J_{\mathrm{12}}/2\pi=62.5$ MHz and hybrid linewidth $(\gamma_{1}+\gamma_{2})/4\pi=3.5$ MHz. c, Calculated eigenstate energy diagram of the three-qubit coupled system, with bare qubit states shown as dashed lines. See Methods for detailed qubit parameters used in the calculation.
Following the frequency-domain characterization of the coupled system, we perform time-domain two-qubit operations corresponding to the CR and bSWAP gates. A Gaussian-shaped microwave pulse, with variable frequency and duration, excites the system and is followed by a readout pulse at the resonator frequency. This readout pulse also functions as the $\mathrm{Q_{b}}$ frequency tuning knob with ac-Stark effect and induces population exchange between $\mathrm{Q_{d}}$ and $\mathrm{Q_{b}}$. The CR and bSWAP gate operations are manifested as oscillations in the readout signal in Fig. Coherent manipulation of interacting electron qubits on solid neona, at 5.726 GHz and 5.718 GHz respectively. The simulation of this dynamic process with the system Hamiltonian gives post-readout population of $\mathrm{Q_{b}}$ with similar patterns, as shown in Fig. Coherent manipulation of interacting electron qubits on solid neonb. The detailed evolution of the two-qubit system through this drive-readout process is presented in Extended Data Fig. 3.
The simulated results match well with the experimental observations. For the CR operation (purple arrows), $\mathrm{Q_{d}}$ was driven by pumping $\mathrm{Q_{b}}$ at the transition frequency of $\mathrm{Q_{d}}$, whose excited state population reached maximum when the pulse length was approximately 0.8 µs. Meanwhile, the CR operation barely excites $\mathrm{Q_{b}}$, as shown in Extended Data Fig. 3. For the bSWAP operation (magenta arrows), the two-photon process drives the two qubits simultaneously, while the oscillation frequency is lower than for the CR operation. When it performs a $\pi$ rotation, which is about 2 µs in our case, the bSWAP gate is equivalent to an iSWAP gate with appropriate single-qubit operations [poletto2012entanglement].
Furthermore, we performed amplitude-dependent Rabi oscillation measurement with a Gaussian-shaped drive pulse of 0.8 µs duration. Again, we observe oscillating patterns near 5.726 GHz and 5.718 GHz, corresponding to the CR (purple arrows) and bSWAP (magenta arrows) operations. We attribute the “waterfall-like” patterns in the frequency range 5.713 through 5.718 GHz in Fig. Coherent manipulation of interacting electron qubits on solid neonc to the effects of the Gaussian shape of the drive pulse and the ac-Stark shift of $\mathrm{Q_{b}}$ under higher drive power, which are reproduced by simulations in Fig. Coherent manipulation of interacting electron qubits on solid neond and Extended Data Fig. 3.
These results demonstrate, for the first time, time-domain two-qubit operations in the eNe qubit platform. The neon surface confinement assisted the global all-microwave gate with simplified local DC bias, offering the potential for compact multi-qubit integration on solid neon. Further optimization of the two-qubit gates and readout is crucial to achieving high-fidelity operations.
Three-qubit strong coupling
We further realized a three-qubit system on solid neon. The experiment was performed during a separate cooldown using a different resonator from that used in the previous two-qubit system. In this three-qubit system, $\mathrm{Q}_{1}$ is directly strong coupled with the resonator, manifested by vacuum Rabi splitting when they are on resonance, while there is no detectable coupling between the other two qubits with the resonator. The energy diagram of the coupled system was probed via a CW two-tone qubit spectroscopy, while varying the gate voltage on a nearby resonator electrode ($\Delta V_{\rm res}$, see Methods). The interactions between qubits are revealed by the avoided crossings in the system spectrum, as shown in Fig. Coherent manipulation of interacting electron qubits on solid neona. The line cut along the bias point where $\mathrm{Q}_{1}$ and $\mathrm{Q}_{2}$ are on resonance, enlarged in Fig. Coherent manipulation of interacting electron qubits on solid neonb, shows a hybridize-induced avoided crossing with a splitting of $J_{\textrm{12}}/2\pi$ = 62.5 MHz. In the same spirit, the coupling strength between $\mathrm{Q}_{2}$ and $\mathrm{Q}_{3}$ is approximately $J_{\textrm{23}}/2\pi$ = 5 MHz. The difference in inter-qubit coupling strengths arises from variations in the distance and dipole orientation between qubits. Given that the couplings between $\mathrm{Q}_{2}$ and $\mathrm{Q}_{3}$ with the resonator are not detectable, likely because of an unfavorable alignment between their dipoles and the microwave field, we attribute the inter-qubit strong coupling to direct charge-charge interactions instead of virtual photon exchange via the resonator.
Figure Coherent manipulation of interacting electron qubits on solid neonc shows the calculated eigenstates of the system in response to the bias voltage applied on the resonator electrode, aligning with the experimental two-tone spectroscopy. The parameters applied in the calculation are listed in Extended Data Tab. Coherent manipulation of interacting electron qubits on solid neon. Due to the large coupling strength between $\mathrm{Q}_{1}$ and $\mathrm{Q}_{2}$, the state mixing of the three qubits results in the energy diagram of the coupled system, which deviates from the uncoupled qubit spectrum. As a result, the three mixed states are all visible in the two-tone measurement in Fig. Coherent manipulation of interacting electron qubits on solid neona. These observations indicate the possibilities of simultaneously entangling more than two qubits in the eNe system that may be applied for small- to medium-scale quantum simulations [wang2023atomic, hensgens2017quantum].
Discussion and outlook
In this work, we have realized coherent manipulation of multiple electron qubits on solid neon mediated by charge-charge interactions. In the two-qubit system, frequency-domain spectroscopy shows strong inter-qubit coupling, while time-domain measurements demonstrate both cross-resonance and bSWAP two-qubit operations, paving the way for the optimizing high-fidelity two-qubit gates on solid neon. The bright qubit in the system act as the mediator between the dark qubit and the control and readout circuitry, showcasing the engineering potential of neon structures for functioning and scaling eNe qubits with compacted integration and suppressed wiring overheads. Further, the three-qubit system with over 60 MHz inter-qubit coupling strength shows the potential of building more complex systems with qubits on solid neon.
Charge-charge interactions between qubits at short distances can enable new possibilities for quantum information processing on solid neon. For the two systems studied in this work, qubits can acquire distinct coupling strengths with superconducting resonators. Leveraging such possibilities, qubits isolated from the resonator can serve as quantum memories for nearby coupled qubits [wendin2017quantum, xie2024high], or act as a quantum mediator [baart2017coherent] in a qubit array. Enabling these applications requires further investigation into engineering the electron’s structural confinement with solid neon to control qubit properties and their interactions.
[…]
Methods
Devices and experiments
The data presented in Fig. Coherent manipulation of interacting electron qubits on solid neon and Fig. Coherent manipulation of interacting electron qubits on solid neon were collected on a high-impedance titanium nitride (TiN) splitting superconducting resonator, the same as the one in ref. [li2025noise]. A schematic of the electron trap on the resonator is shown in Extended Data Fig. 1a. The resonator supports a differential mode at 5.668 GHz with a total linewidth of $\kappa/2\pi$ = 0.38 MHz. Electrodes with on-chip low-pass filters are connected to the resonator pins and trap guards surrounding the electron-trapping region, enabling the application of bias voltages to tune the qubit transition frequency. The device was mounted on a home-made printed circuit board (PCB), sealed in a copper cell with neon filling lines, tungsten filaments, and electrical connectors on top of the cell lid. The cryogenic and room temperature electronics setups are the same as those in ref. [li2025noise]. The data presented in Fig. Coherent manipulation of interacting electron qubits on solid neon were collected on a niobium splitting superconducting resonator [zhou2022single, zhou2024electron], which supports a differential mode at 6.426 GHz, with a total linewidth of $\kappa/2\pi$ = 0.46 MHz. The sample, cryogenic and room temperature electronic setups were the same as those in refs. [zhou2022single, zhou2024electron].
The neon growth procedure is the same as that described in ref. [li2025noise]. In summary, the dilution fridge is warmed up to create a temperature gradient from $\sim$27 K to $\sim$25 K between its 4K plate and mixing chamber (MXC) plate. At that moment, neon gas is supplied from a room-temperature gas handling system in the form of puffs and filled onto the device chip in the form of liquid. After a desired amount of neon filling, the fridge heater is turned off to let the fridge cool down to the base temperature.
Electrons are deposited onto the chip at the base temperature by applying pulsed current to the tungsten filaments mounted above the device. During the electron firing process, temperature sensed on the MXC plate can increase up to 100 mK. The details of the procedure and discussion of the electron deposition process can be found in ref. [li2025noise].
Modeling multi-qubit coupled systems
The multi-qubit system Hamiltonian can be written as
$\hat{H}_{\mathrm{sys}}=\hat{H}_{\mathrm{r}}+\hat{H}_{\mathrm{q}}+\hat{H}_{\mathrm{r\text{-}q}}+\hat{H}_{\mathrm{q\text{-}q}},$ (M1)